Shift report 20000830 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Wed Aug 30 2000 - 08:31:03 EDT

  • Next message: http account: "Shift report 20000830 08:00-16:00"

    ------------------------------------------------------------
    Date: Wed Aug 30 08:31:02 2000, Shift: 00:00-08:00
    Supervisor: FR, with: 
    Log book#5 pages 49 - 51
    
    Shift Summary:
    ------------------------------------------------------------
    
     1) Continuous beam since 0:55. Intensities were b=14900 and 
     y=15100. The beam was stable all night with less background
     than in the previous store (vacuum is improving).
     There was a problem with trigger 4 due to yellow clock 
     timing. DB moved yellow clock to previous value (see Log
     book, p.51).
    
     2) Three data files recorded with a total of 1021K 
     trigger-1 events. 
     Run 2472 has 326K trigger-1 events and about 177K coll.
     Run 2473 has 391K trigger-1 events and about 204K coll.
     Run 2474 has 304K trigger-1 events and about 160K coll.  
     All above runs have been taken with MRS at 90deg., FFS 
     at 5deg. and B polarity for D1/D2/D5.
    
     3) No detector problems observed.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Wed Aug 30 2000 - 08:34:42 EDT